# Squared-field cross-correlation between kinetic Sunyaev-Zel’dovich effect and 21-cm intensity mapping

### v1

@@ IV Conclusions and discussions
In this paper, we established the analytical formulae for the cross-correlation power spectrum between the kSZ square field and the projection of HI square field under the Gaussian and flat-sky approximations. In terms of the validity of the approximation, for the low-redshifts HI intensity mapping field ($z\lesssim 3$) where reionizations of hydrogen and helium are fully completed, the Gaussian approximation works well for the perturbed HI intensity field that follows the matter density field. However, we caution that these formulae may be inaccurate at higher redshifts during reionization ($6\lesssim z\lesssim 15$), where the ionized regions lead to a strongly non-Gaussian signal while spin temperature fluctuations may lead to further complications, at least during early phases of the EoR Mesinger et al. (2011). In that scenario, the analytical formula we can achieve at most is Eq. (24), in which multi-point correlation $\mathcal{T}$ must be evaluated in a fully numerical way. For the flat-sky approximation, it is reasonable as long as we consider high-$\ell$ modes of correlation, in which the curvature of the sky can be safely neglected. To calculate the complete, full-sky power spectrum, one needs to work with the spherical-harmonics decomposition instead of 2D plane-wave Fourier transformation, which is left to further study.
To demonstrate the results, we numerically calculated the cross-correlation power spectrum in the range of $30\leq\ell\leq 900$, between the redshift $0.3\leq z\leq 1$ which is the sensitive range of SKA-MID IM project. We forecast that the SNR for SKA-MID to reach a maximum value $\mathrm{SNR}_{\mathrm{max}}\simeq 1.92$ for ACT and $\mathrm{SNR}_{\mathrm{max}}\simeq 3.99$ for SO. These results come from the assumption of constant biases for the electron and HI density contrasts, which is reasonable because the main contributions in Eq. (41) come from the peak of the power spectrum, where the non-linear effects are safely neglected. Future investigations can be done to calculate the scale-dependent biases derived from the halo model.
One of the advantages of taking square field cross-correlation is that the convolution arising from square field helps to leverage the $k^{\parallel}$ modes in 22D map cross-correlations, catching more information about the large-scale structure of the universe than simple projection fields. The $k^{\parallel}$ modes will become more important when the evolution of the structure is not negligible, for example, during the EoR period. Future studies can be carried out by evaluating the $\mathrm{kSZ}^{2}\times\mathrm{HI}^{2}$ cross-correlation during EoR for HI signal detection and exploring the physics of the first stars and galaxies at high redshifts (e.g. Ref. Zhou et al. (2025)).
@@
As a final remark, the cross-correlations between HI intensity mapping and other tracers of large-scale structure can serve as a validation of methodology for high-redshift 21-cm detection, which faces severe challenges from foreground contamination, especially when coupled with instrumental effects (e.g. mutual coupling effect Charles et al. (2024); Rath et al. (2025) and foreground leakage Orosz et al. (2019); Gogo et al. (2022)). The kSZ signal during the EoR is also one of the few accessible signals across large regions of the sky. At low redshifts, we carried out a detailed study on cross-correlation of kSZ square field and HI square field, which provides a presumably easier test of the statistics. A validation of this method at low redshifts can provide further motivation to pursue similar procedure back into the EoR. In addition, at lower redshifts, the expected kSZ signals are better understood, and there are already detected cross-correlations with galaxy surveys Hand et al. (2012); Li et al. (2018); Ade et al. (2016) and velocity reconstruction fields Ade et al. (2016); Hernández-Monteagudo et al. (2015). Our method opened a new window by novelly introducing their cross-correlation with 21-cm signal, therefore provides a testing ground and cross-check for other detection statistics.
@@ Acknowledgements.
The authors acknowledge the useful discussions with Meng Zhou, Steven Murray and the MeerKLASS collaboration team. Z.-Y. Yuwen and Y.-E. Jiang are supported by the Program of China Scholarship Council Grant No. 202404910329 and No. 202404910398. Y.-Z. Ma acknowledges the support from South Africa’s National Research Foundation under grant Nos. 150580, CHN22111069370, ERC250324306141. P. L. Plante is supported by NSF Grant No. 2206602 and the Simons Foundation award No. 00007127. All computations were carried out using the computational cluster resources at the Centre for High-Performance Computing, Cape Town, South Africa.
@@
For the SP field, one can first get the projection of HI signal along line-of-sight given by Eq. (5) and then square it, from which the SP field reads
@@
In comparison, replacing SP field with PS field suggests a different cross-correlation
@@
We now apply Limber approximation, which assumes a slow-varying parallel modes dependence in the bispectrum for small-angular scale satisfying $k^{\parallel}\ll\ell/\chi$ (as long as the mode is not enrolled in the convolution). This allows us to neglect the $k^{\parallel}$ in $\mathcal{B}^{\mathrm{HI}^{2}\mathrm{g}}$ and simplify Eqs. (A) and (A) to the following expressions
@@
where $\cos\theta=\hat{k}\cdot\hat{\ell}$. The last line is derived by noticing that the bispectrum should be a function of $|\vec{k}|$, $\ell/\chi$ and the separation angle $\cos\theta$. Similar result remains correct if one replaces $\delta_{\mathrm{g}}$ with other projected field such as kSZ, as is given in Eq. (35).
One can clearly see the difference between Eqs. (66) and (67), which is the momentum integral of the bispectrum. The PS method results in a 3-D integral which fully takes the advantage of parallel momentum mode, whereas SP method loses this $k^{\parallel}$ information and only utilizes the $k^{\perp}$ modes. Therefore, the $C_{\ell}^{\mathrm{PS}\times{\mathrm{g}}}$ has significantly higher signal than $C_{\ell}^{\mathrm{SP}\times{\mathrm{g}}}$.
Appendix B Redshift-space distortion effect of HI intensity mapping
@@
Figure 6: Correlation functions with and without RSD effect. Left–Power spectrum of HI square field; Right–Each individual term contribution in the cross-power spectrum between ${\rm kSZ}^{2}$ and ${\rm HI}^{2}$ fields. The absolute value of total cross-correlation $C_{\ell}^{\mathrm{tot}}$ coincides with $C_{\ell}^{(6)}$ because of the hierarchies among contributions from different terms.
In this appendix, we discuss how the redshift-space distortion (RSD) correction on HI intensity mapping is applied to the density contrasts (and further to the correlations functions). In a realistic HI intensity mapping observation, the observable is a radiative quantity—the brightness $I_{\nu}$, which is defined by the received energy $\mathrm{d}E$ per unit time $\mathrm{d}t$ per detector area $\mathrm{d}A$ per solid angle $\mathrm{d}\Omega$ within the frequency range $(\nu,\nu+\mathrm{d}\nu)$
@@
where $\mathcal{H}=aH$ is the comoving Hubble parameter. We have neglected the relatively small local evolution of the background, lensing and general-relativistic effects, which more details can be found in Ref. Hall et al. (2013). The energy conservation then reads
@@
where the subscript $s$ denotes for quantities in redshift space. In the Rayleigh-Jeans regime, the brightness temperature is proportional to the $I_{\nu}$ function, i.e. $T_{\mathrm{b}}\simeq I_{\nu}c^{2}/(2k_{\mathrm{B}}\nu^{2})$, one can then rewrite Eq. (71) as
@@
Here we have used the fact that perturbation of brightness temperature $\delta_{T_{\mathrm{b}}}$ equals the density contrast of HI $\delta_{\mathrm{HI}}$. Further using the relation (70) results in
@@
While evaluating the power spectra, the presence of RSD effect means the replacement: $\delta_{\mathrm{HI}}\to\delta_{s,\mathrm{HI}}$. Then the HI field auto correlation power spectrum now becomes
@@
which is nothing but the classical Kaiser effect Kaiser (1984). The same replacement can be applied to other correlation functions. The inclusion of RSD effect in auto and cross-correlation power spectra are shown in Fig. 6. Similar to HI auto-correlation case, the amplitude of the power spectrum is increased about $1.5$ times with the inclusion of RSD effect, while the shape of it remains not affected.
@@ Appendix C Cross-correlation power spectrum in the fine-beam limit $\theta^{\rm HI}_{\rm FWHM}\rightarrow 0$
Figure 7: Cross-correlation angular power spectrum for HI IM in the redshift range $0.3<z<0.6$ in fine-beam limit ($\theta^{\rm HI}_{\rm FWHM}\rightarrow 0$), with biases $b_{\mathrm{e}}=b_{\mathrm{HI}}=1$.
It is obvious that the cross-correlation power spectrum (Eq. (41)) is suppressed by both kSZ filter and HI beam. The former is hard to modify on large scales because it comes from the ratio of power between kSZ effect and primary CMB, which is order of $10^{-8}$ in general. However, the latter can be improved by enlarging the observation baseline, which can be achieved by using a much larger dish such as FAST telescope with effective diameter $D=300\mathrm{m}$ Nan et al. (2011), or the interferometry experiments with longer baselines Bacon et al. (2020). Here we briefly discuss the fine-beam limit of HI ($\theta^{\rm HI}_{\rm FWHM}\rightarrow 0$, i.e. $B_{\mathrm{HI}}\to 1$) to see the capacity of signal enhancement.
We first recall that $Y^{(7)}$ given by Eq. (II.3) contains $P_{v,v}(k)$ factor, which has a “resonant divergence” as $k\to 0$, corresponding to the condition $\vec{l}_{1}+\vec{p}\chi\to 0$. For a rough estimation, one can neglect the change of the other two $P_{\mathrm{e,HI}}$ factors in the vicinity of the pole $\vec{l}_{1}+\vec{p}=0$. However, such a contribution from the pole will be highly suppressed at large $p^{\perp}=\ell_{1}/\chi$ with $p^{\parallel}=0$, since it has been killed by HI beam. Fine-beam limit allows high $p_{\perp}$ to survive, making contribution to an effective factor proportional to root-mean-square of velocity perturbation. Then the integral over $\vec{p}$ can be written as
@@
Here we show the numerical results of the cross-correlation in Fig. 7 within the redshift window $0.3<z<0.6$. From Fig. 7, we can see that although $C^{(6)}_{\ell}$ still dominate the contribution on large scales, $C^{(7)}_{\ell}$ gradually takes over as $\ell$ grows and eventually reaches a peak at $\ell\simeq 6000$, and then decays at smaller scales because of kSZ filter. In this numerical calculation we did not include foreground filter to reach the limit $k^{\parallel}\to 0$. By comparing Fig. 7 with Fig. 2, one can see that Term 7 and the total power spectrum can be enhanced to $\mathcal{O}(5\times 10^{-1})$ if $\ell\simeq 6000$ can be measured. However, this indicates that the dish size or interferometer baseline has to be at least $\sim 605$ meter [^6] We calculate the dish size by equating $\theta=\frac{\lambda}{D}=\frac{\pi}{\ell}\Rightarrow D=\frac{\lambda\ell}{\pi}=\left(\frac{c}{\nu}\right)\frac{\ell}{\pi}\simeq 605\,{\rm m}.$ (78) , which excludes the possibility of single-dish measurement, but resorts to the interferometer measurements.

### v2

@@ IV Conclusions and discussions
In this paper, we established the analytical formulas for the cross-correlation power spectrum between the kSZ square field and the projection of the HI square field under the Gaussian and flat-sky approximations. In terms of the validity of the approximation, for the low-redshifts HI intensity mapping field ($z\lesssim 3$) where reionizations of hydrogen and helium are fully completed, the Gaussian approximation works well for the perturbed HI intensity field that follows the matter density field. However, we caution that these formulas may be inaccurate at higher redshifts during reionization ($6\lesssim z\lesssim 15$), where the ionized regions lead to a strongly non-Gaussian signal while spin temperature fluctuations may lead to further complications, at least during early phases of the EoR Mesinger et al. (2011). In that scenario, the analytical formula we can achieve at most is Eq. (24), in which multipoint correlation $\mathcal{T}$ must be evaluated in a fully numerical way. For the flat-sky approximation, it is reasonable as long as we consider high-$\ell$ modes of correlation, in which the curvature of the sky can be safely neglected. To calculate the complete, full-sky power spectrum, one needs to work with the spherical-harmonics decomposition instead of 2D plane-wave Fourier transformation, which is left to further study.
To demonstrate the results, we numerically calculated the cross-correlation power spectrum in the range of $30\leq\ell\leq 900$, between the redshift $0.3\leq z\leq 1$ which is the sensitive range of the SKA-MID IM project. We forecast the SNR for SKA-MID to reach a maximum value $\mathrm{SNR}_{\mathrm{max}}\simeq 1.92$ for ACT and $\mathrm{SNR}_{\mathrm{max}}\simeq 3.99$ for SO. These results come from the assumption of constant biases for the electron and HI density contrasts, which is reasonable because the main contributions in Eq. (41) come from the peak of the power spectrum, where the nonlinear effects are safely neglected. Future investigations can be done to calculate the scale-dependent biases derived from the halo model.
One of the advantages of taking the square field cross-correlation is that the convolution arising from the square field helps to leverage the $k^{\parallel}$ modes in 22D map cross-correlations, catching more information about the large-scale structure of the universe than simple projection fields. The $k^{\parallel}$ modes will become more important when the evolution of the structure is not negligible, for example, during the EoR period. Future studies can be carried out by evaluating the $\mathrm{kSZ}^{2}\times\mathrm{HI}^{2}$ cross-correlation during EoR for HI signal detection and exploring the physics of the first stars and galaxies at high redshifts (e.g. Ref. Zhou et al. (2025)).
@@
As a final remark, the cross-correlations between HI intensity mapping and other tracers of large-scale structure can serve as a validation of methodology for high-redshift 21-cm detection, which faces severe challenges from foreground contamination, especially when coupled with instrumental effects (e.g. the mutual coupling effect Charles et al. (2024); Rath et al. (2025) and foreground leakage Orosz et al. (2019); Gogo et al. (2022)). The kSZ signal during the EoR is also one of the few accessible signals across large regions of the sky. At low redshifts, we carried out a detailed study on the cross-correlation of the kSZ square field and HI square field, which provides a presumably easier test of the statistics. A validation of this method at low redshifts can provide further motivation to pursue similar procedure back into the EoR. In addition, at lower redshifts, the expected kSZ signals are better understood, and there are already detected cross-correlations with galaxy surveys Hand et al. (2012); Li et al. (2018); Ade et al. (2016) and velocity reconstruction fields Ade et al. (2016); Hernández-Monteagudo et al. (2015). Our method opened a new window by novelly introducing their cross-correlation with a 21-cm signal, providing a testing ground and cross-check for other detection statistics.
@@ Acknowledgements.
The authors acknowledge the useful discussions with Meng Zhou, Steven Murray and the MeerKLASS collaboration team. Z.-Y. Yuwen and Y.-E. Jiang are supported by the Program of China Scholarship Council Grant No. 202404910329 and No. 202404910398. Y.-Z. Ma acknowledges the support from South Africa’s National Research Foundation under Grant No. 150580, No. CHN22111069370, No. ERC250324306141. P. L. Plante is supported by NSF Grant No. 2206602 and the Simons Foundation award No. 00007127. All computations were carried out using the computational cluster resources at the Centre for High-Performance Computing, Cape Town, South Africa.
@@
For the SP field, one can first get the projection of the HI signal along the line-of-sight given by Eq. (5) and then square it, from which the SP field reads
@@
In comparison, replacing the SP field with the PS field suggests a different cross-correlation
@@
We now apply Limber approximation, which assumes a slow-varying parallel modes dependence in the bispectrum for a small-angular scale satisfying $k^{\parallel}\ll\ell/\chi$ (as long as the mode is not enrolled in the convolution). This allows us to neglect the $k^{\parallel}$ in $\mathcal{B}^{\mathrm{HI}^{2}\mathrm{g}}$ and simplify Eqs. (A) and (A) to the following expressions
@@
where $\cos\theta=\hat{k}\cdot\hat{\ell}$. The last line is derived by noticing that the bispectrum should be a function of $|\vec{k}|$, $\ell/\chi$ and the separation angle $\cos\theta$. A similar result remains correct if one replaces $\delta_{\mathrm{g}}$ with another projected field such as kSZ, as is given in Eq. (35).
One can clearly see the difference between Eqs. (66) and (67), which is the momentum integral of the bispectrum. The PS method results in a 3D integral which fully takes the advantage of parallel momentum mode, whereas SP method loses this $k^{\parallel}$ information and only utilizes the $k^{\perp}$ modes. Therefore, the $C_{\ell}^{\mathrm{PS}\times{\mathrm{g}}}$ has significantly higher signal than $C_{\ell}^{\mathrm{SP}\times{\mathrm{g}}}$.
Appendix B Redshift-space distortion effect of the HI intensity mapping
@@
Figure 6: Correlation functions with and without RSD effect. Left–Power spectrum of the HI square field; Right–Each individual term contribution in the cross-power spectrum between ${\rm kSZ}^{2}$ and ${\rm HI}^{2}$ fields. The absolute value of total cross-correlation $C_{\ell}^{\mathrm{tot}}$ coincides with $C_{\ell}^{(6)}$ because of the hierarchies among contributions from different terms.
In this appendix, we discuss how the redshift-space distortion (RSD) correction on HI intensity mapping is applied to the density contrasts (and further to the correlations functions). In a realistic HI intensity mapping observation, the observable is a radiative quantity – the brightness $I_{\nu}$, which is defined by the received energy $\mathrm{d}E$ per unit time $\mathrm{d}t$ per detector area $\mathrm{d}A$ per solid angle $\mathrm{d}\Omega$ within the frequency range $(\nu,\nu+\mathrm{d}\nu)$
@@
where $\mathcal{H}=aH$ is the comoving Hubble parameter. We have neglected the relatively small local evolution of the background, lensing and general-relativistic effects, for which more details can be found in Ref. Hall et al. (2013). The energy conservation then reads
@@
where the subscript $s$ denotes quantities in redshift space. In the Rayleigh-Jeans regime, the brightness temperature is proportional to the $I_{\nu}$ function, i.e. $T_{\mathrm{b}}\simeq I_{\nu}c^{2}/(2k_{\mathrm{B}}\nu^{2})$, and one can then rewrite Eq. (71) as
@@
Here we have used the fact that the perturbation of brightness temperature $\delta_{T_{\mathrm{b}}}$ equals the density contrast of the HI $\delta_{\mathrm{HI}}$. Further using the relation (70) results in
@@
While evaluating the power spectra, the presence of the RSD effect means the replacement: $\delta_{\mathrm{HI}}\to\delta_{s,\mathrm{HI}}$. Then the HI field auto correlation power spectrum now becomes
@@
which is nothing but the classical Kaiser effect Kaiser (1984). The same replacement can be applied to other correlation functions. The inclusion of RSD effect in auto- and cross-correlation power spectra are shown in Fig. 6. Similar to HI autocorrelation case, the amplitude of the power spectrum is increased about $1.5$ times with the inclusion of the RSD effect, while the shape of it remains not affected.
@@ Appendix C Cross-correlation power spectrum in the fine-beam limit $\theta^{\rm HI}_{\rm FWHM}\rightarrow 0$
Figure 7: Cross-correlation angular power spectrum for HI IM in the redshift range $0.3<z<0.6$ in the fine-beam limit ($\theta^{\rm HI}_{\rm FWHM}\rightarrow 0$), with biases $b_{\mathrm{e}}=b_{\mathrm{HI}}=1$.
It is obvious that the cross-correlation power spectrum [Eq. (41)] is suppressed by both the kSZ filter and the HI beam. The former is hard to modify on large scales because it comes from the ratio of power between the kSZ effect and primary CMB, which is order of $10^{-8}$ in general. However, the latter can be improved by enlarging the observation baseline, which can be achieved by using a much larger dish such as FAST telescope with effective diameter $D=300\mathrm{m}$ Nan et al. (2011), or the interferometry experiments with longer baselines Bacon et al. (2020). Here we briefly discuss the fine-beam limit of the HI ($\theta^{\rm HI}_{\rm FWHM}\rightarrow 0$, i.e. $B_{\mathrm{HI}}\to 1$) to see the capacity of signal enhancement.
We first recall that $Y^{(7)}$ given by Eq. (II.3) contains the $P_{v,v}(k)$ factor, which has a “resonant divergence” as $k\to 0$, corresponding to the condition $\vec{l}_{1}+\vec{p}\chi\to 0$. For a rough estimation, one can neglect the change of the other two $P_{\mathrm{e,HI}}$ factors in the vicinity of the pole $\vec{l}_{1}+\vec{p}=0$. However, such a contribution from the pole will be highly suppressed at large $p^{\perp}=\ell_{1}/\chi$ with $p^{\parallel}=0$, since it has been killed by the HI beam. The fine-beam limit allows high $p_{\perp}$ to survive, making a contribution to an effective factor proportional to the root-mean-square of velocity perturbation. Then the integral over $\vec{p}$ can be written as
@@
Here we show the numerical results of the cross-correlation in Fig. 7 within the redshift window $0.3<z<0.6$. From Fig. 7, we can see that although $C^{(6)}_{\ell}$ still dominate the contribution on large scales, $C^{(7)}_{\ell}$ gradually takes over as $\ell$ grows and eventually reaches a peak at $\ell\simeq 6000$, and then decays at smaller scales because of the kSZ filter. In this numerical calculation we did not include foreground filter to reach the limit $k^{\parallel}\to 0$. By comparing Fig. 7 with Fig. 2, one can see that Term 7 and the total power spectrum can be enhanced to $\mathcal{O}(5\times 10^{-1})$ if $\ell\simeq 6000$ can be measured. However, this indicates that the dish size or interferometer baseline has to be at least $\sim 605$ m [^6] We calculate the dish size by equating $\theta=\frac{\lambda}{D}=\frac{\pi}{\ell}\Rightarrow D=\frac{\lambda\ell}{\pi}=\left(\frac{c}{\nu}\right)\frac{\ell}{\pi}\simeq 605\,{\rm m}.$ (78) , which excludes the possibility of single-dish measurement, but resorts to the interferometer measurements.